OpenStudy (anonymous):

Given polygon BRAIN is similar to polygon POWEQ Calculate the value of WE, BR, and EQ A. WE = 5, BR = 11.25, EQ = 7 B. WE = 11.25, BR = 5, EQ = 7 C. WE = 5, BR = 7, EQ = 11.25 D. WE = 7, BR = 11.25, EQ = 5

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

i just need help getting the equations the rest i can do

OpenStudy (aripotta):

WE- 10/4 = 12.5/(2x + 1) BR- 10/4 = (y + 5)/4.5 EQ- 10/4 = 17.5/14z

OpenStudy (anonymous):

if u got time can u show me how u got them please